### v2

# Improved and Explainable Cervical Cancer Classification using Ensemble Pooling of Block Fused Descriptors

Abstract
Cervical cancer is the second most common cancer in women and causes high death rates. Earlier models for detecting cervical cancer had limited success. In this work, we propose new models that substantially outperform previous models.
Previous studies show that pretrained ResNets extract features from cervical cancer images well. Hence, our first model involves working with three ResNets (50, 101, 152). All the existing works use only the last convolution block of their respective ResNet, which captures abstract features (e.g., shapes, objects). However, we believe that detailed features (e.g., color, edges, texture), coming from earlier convolution blocks, are equally important for cancer (specifically cervical cancer) classification. Since now the number of features become large, we use a novel feature selection technique of Global Max Pooling for detailed features and Global Average Pooling for abstract features. Hence, our second model consists of the resulting Cascaded Block Fused variants of the three ResNets. To improve the performance further, we combine and normalize the features of the three standard ResNets as well as our proposed three Cascaded Block Fused ResNets. This type of combination is also new in cancer classification domain (also in cervical cancer), and results in our third and fourth models, respectively. We use a linear SVM for classification.
We exhaustively perform experiments on two public datasets, IARC and AnnoCerv, achieving an average performance of $97.92\%$ and $92.97\%$ surpassing standard ResNets with performance of 90.89% and 87.97%, respectively. We outperform the competitive approach available on IARC dataset with an average gain of $13.20\%$, while no prior competitive work available on AnnoCerv. Additionally, we introduce a novel SHAP+LIME explainability method, accurately identifying the cancerous region in 97% of cases, ensuring model reliability for real-world use.
keywords:
Cervical cancer, Colposcopy, Pretrained Convolution Neural Networks, ResNet, , Global Max Pooling, Global Average Pooling, Descriptor Hybridization, Support Vector Machine, Explainable AI, SHAP, LIME
1 Introduction
Cervical cancer is a serious health concern for women aged 15–44, ranking second in fatality after breast cancer with 340,000 deaths annually [1]. It is also the second most diagnosed cancer in women, with more than 600,000 annual cases [2]. The main cause of this is the Human Papilloma Virus (HPV) transmitted through sexual contact [3].
Early screening programs can save lives since this cancer is considered incurable in the advanced stages. Screening methods include visual inspection with acetic acid (VIA), HPV test, pap smear, and colposcopy [1]. Colposcopy is a widely used method that involves taking pictures of the cervix after applying acetic acid [4]. This helps identify the various degrees of abnormalities visible in shades of white [5]. There is a shortage of expert healthcare professionals who can accurately classify colposcopy cancer images (into different types and Cervical Intraepithelial Neoplasia (CIN) scores). Hence, Computer Aided Diagnosis (CAD) systems are preferred. Developing a CAD system for colposcopy cancer image classification that works substantially better than the best in literature forms the focus of this work. A typical CAD system consists of three primary components, namely, feature extraction, classification, and explainability.
Feature extraction can be performed using either handcrafted descriptors or deep learning descriptors [6]. Handcrafted descriptors do not capture all the features, while the deep learning descriptors capture both the detailed features (e.g., edges, color, and texture) as well as the abstract features (e.g., shapes and objects) well [7, 8]. For colposcopy cancer image classification, all kinds of features are important [9], and hence, this makes deep learning a better choice here.
In deep learning, Convolutional Neural Networks (CNNs) are a popular choice. They typically come with a variety of pretrained models, such as Inception, VGG16, ResNet, AlexNet, etc. [10], which are highly reliable because they are trained on large datasets like ImageNet. In the field of medical imaging, it has been recently shown that ResNets are more accurate than the other pretrained neural networks [11, 12, 13, 14]. Further, one of the common problems in deep neural networks is the vanishing gradient problem, which these ResNets also address. Hence, in this work, ResNets form our underlying pretrained neural network. Since lower-numbered ResNets capture simpler patterns while the higher-numbered ResNets capture more complex patterns, we extract features using ResNet50, ResNet101, and ResNet152, which forms our first model.
It has been shown in pretrained neural networks that earlier convolution blocks capture detailed features such as color, edges, and texture, while the later convolution blocks capture abstract features like shapes and objects. These abstract features dominate the output of the network. For colposcopy cancer image classification, detailed features are equally important. The significance of detailed features was first highlighted in [15], which laid the foundation for their use in various applications. Following this, several studies, including [16] and [17], applied this concept to the task of scene classification. Despite their success, this idea of combining these types of features has not been explored in the context of any type of cancer classification, and certainly not in colposcopy cancer image classification, which has led the previous studies to the relatively low performance [18, 12, 13, 19, 1]. To address this issue, we extract features at the end of each convolution block in the three selected ResNets (ResNet50, ResNet101, and ResNet152) in a block-wise cascading manner. This gives us both detailed features and abstract features that lead to highly accurate colposcopy cancer classification.
Since the feature space become very large, hence, we need to apply feature selection. There are several ways to select features in CNNs, such as Global Sum Pooling (GSP), Global Max Pooling (GMP), and Global Average Pooling (GAP) [20, 21]. As discussed earlier, the initial convolution blocks in CNNs primarily capture detailed features, which are often sensitive to background noise. GMP is known for its ability to suppress irrelevant patterns [22], therefore, we apply GMP to the detailed features extracted from the first two convolution blocks of each selected ResNet. On the other hand, the later convolution blocks extract abstract features that are more focused on the shape and structure of the object and are not affected by background noise. Thus, applying GAP is appropriate for these layers, and we use it on the features from the last three convolution blocks of each ResNet [22]. For each ResNet, the output from all the GMP and GAP layers is fused together. As a result, we obtain our new Cascaded Block Fused ResNet descriptors, i.e., CBF-ResNet50, CBF-ResNet101, and CBF-ResNet152, corresponding to ResNet50, ResNet101, and ResNet152, respectively. This set constitutes our second model.
In multiple domains such as clustering and classification, it has been shown that combining features from several pre-trained CNNs helps to improve the performance [23, 24]. Following this approach, we combine the and normalize features of the three standard ResNets (ResNet50, ResNet101, and ResNet152) and our three newly proposed CBF-ResNets (CBF-ResNet50, CBF-ResNet101, and CBF-ResNet152), resulting in our third and fourth models, respectively.
After feature extraction and selection, we need to classify the colposcopy cancer images. When the two parts are combined together, then the model is referred to as an end-to-end model that typically works well on large to medium-sized datasets. When feature extraction and selection are done separately from classification, then the model is referred to as a hybrid model that works well even on medium to small-sized datasets [6]. The publicly available datasets under consideration here (see below) are medium to small, and hence, we work with a hybrid model while using a linear Support Vector Machine (SVM) for classification. To evaluate our CAD system, we perform both the 5-fold and the 10-fold cross-validation. In each case, the dataset is divided into training and testing sets. The training and testing sets are normalized separately before being fed into the SVM for classification.
There are only two publicly available high-resolution colposcopy cancer image datasets [^1] There is also an older public dataset called the Intel & MobileODT dataset [25], which is available on Kaggle. However, the image quality is low because the images were captured using mobile-level devices [26]., namely, IARC [27] and AnnoCerv [28]. We perform our experiments on both the IARC and the AnnoCerv datasets. Originally, the IARC dataset contains $571$ images, and based upon a recent previous work [18], we apply data augmentation techniques to expand this dataset, resulting in $4725$ images. The standard ResNets obtain an average performance of $90.89\%$, while our best model obtain an average performance of ${\bf 97.92\%}$ on this dataset for the cancer types classification. On the contrary, the best approach in literature only achieved an average performance of $84.46\%$ on the same dataset. The AnnoCerv dataset originally contains $531$ images, which we again augment with the same approach to obtain a total of $4665$ images. On this dataset, the standard pretrained CNNs (ResNets) achieved an average performance of $87.97\%$, while our model achieves an average performance of ${\bf 92.97\%}$ for the CIN score classification, with no competitive approach available for comparison.
Next, we look at explainability. To enhance the trustworthiness and interpretability of the model outputs for healthcare experts, we apply an ensemble of XAI techniques, namely, SHAP (Shapley Additive Explanations) and LIME (Local Interpretable Model-agnostic Explanations), to highlight the region of the colposcopy images that contributes more in the decision-making process. As far as we know, this combination of XAI techniques is being done for the first time.
Cervical cancer usually starts in the cervix, hence, for a model to be trusted by healthcare experts, it should pay attention to that area. The region identified by our ensemble XAI technique on a subset of our first dataset, which contributes to the classification decision, is located around the cervix in ${\bf 97\%}$ of the cases. This shows that the decisions of our model match what healthcare experts expect. This not only proves our classification is correct but also makes the system clearer and more trusted in real medical use.
To summarize, this research offers four distinct contributions:
  • Standard ResNets have demonstrated strong performance in colposcopy cancer image classification. Based on this, we propose new variants of these ResNet that capture both the detailed and the abstract features. Additionally, we introduce a novel feature selection technique to select the features, resulting in our proposed CBF-ResNets. This forms our first two sets of models.
  • We combine and normalize the features of standard ResNets to form our third model and the features of our proposed CBF-ResNets to form our fourth model. We use a linear SVM for classification.
  • We exhaustively test on two publicly available datasets and achieve performance ranging from mid-nineties to the late-nineties for both datasets, which is considered to be excellent. Our results surpass all previously reported works in the literature.
  • Among the available XAI models, it is difficult to identify the one which works best for a particular classification model. Hence, we introduce a new ensemble of two XAI techniques, namely, SHAP and LIME, to explain the classification decisions, which give performance close to hundred percent. This makes our CAD system transparent and clinically trustworthy.
The rest of this manuscript consists of five more sections. In Section 2, we review the existing literature. The methodology is described in Section 3. In Section 4, we discuss the dataset and give the numerical results. In Section 5, we explain the decision of our CAD system using explainable AI techniques. Finally, the conclusion and the future work are given in Section 6.
2 Literature review
Table 1 provides a summary of existing work on colposcopy cancer image classification. These studies performed different types of classification (2-way, 3-way, and 5-way) based on the annotations available provided in the respective datasets.
Xu et al. [1] in $2017$ performed a 2-way (normal-abnormal) classification. They explored both the end-to-end and the hybrid approaches. For the end-to-end approach, they utilized a pretrained CaffeNet model. For hybrid, they used PLAB, PHOG, PLBP, and CaffeNet for feature extraction, and SVM and AdaBoost were used for classification. For this study, the author used a total of $1112$ images from the private dataset. The results demonstrated that the end-to-end deep learning model slightly outperformed the hybrid model, achieving an accuracy rate of $83.42\%$.
Saini et al. [19] in $2020$ also performed a 2-way classification. They used an end-to-end CNN model referred to as ColpoNet. This model was evaluated on a private dataset, which contained a total of $800$ images. They achieved an accuracy of $81.35\%$.
Yan et al. [12] in $2021$ again performed a 2-way classification. They used an end-to-end pretrained CNN model, namely, ResNet18. For this study, a private dataset containing $1400$ image has been used to evaluate the performance of the model. This model also achieved an accuracy of $85.50\%$.
Yuan et al. [13] in $2021$ also performed a 2-way classification. They used an end-to-end pretrained ResNet50 model. To evaluate their model, they used a private dataset of $22330$ images. The model attained an accuracy of $84.10\%$.
Cho et al. [29] in $2020$ designed a model for 5-way classification. This classification was based on Cervical Intraepithelial Neoplasia (CIN) scores. Here, the images are categorized into normal, CIN1, CIN2, CIN3, and cancer. This was an end-to-end model. They used two pre-trained CNNs, namely, Inception-ResNet-V2 and ResNet152. They tested their model on a private dataset consisting of $791$ images. They showed that ResNet152 achieved higher performance than Inception-ResNet-V2, with an accuracy of $51.70\%$.
Recently, Dash et al. [18] in $2023$ conducted a 3-way classification. Here, the dataset is categorized into three types based on their Transformation Zone (TZ), namely, Type1, Type2, and Type3. This was a hybrid model, where they used a modified Inception-ResNet-V2 model to extract the features at multiple scales from the colposcopy images and merge them. They used a linear SVM for classification. They evaluated their model on $292$ images taken from the IARC dataset. Their results showed an accuracy of $81.24\%$.
Most of the above works have worked with standard ResNets. In this work, we propose a family of ResNet models which exploits the fact that the detailed features, which come from the earlier convolution blocks of the ResNet, are equally important as the abstract features. Furthermore, while previous studies have evaluated their models on a single dataset, we assess our models on two publicly available datasets, demonstrating substantial performance gains over existing approaches.
Table 1: Past work in classification of colposcopy images
[…]
tion Type Model Technique tion model
Xu et al. [1] 2-way End-to- CaffeNet, Private data- $83.40$ $88.30$ $83.42$
(2017) classifica- End, set from NCI
tion Hybrid PHOG, SVM, Guanacaste
PLAB, AdaBoost ($1112$ images)
PLBP [30]
Saini et al. [19] 2-way End-to- ColpoNet Private data- $81.35$
(2021) classifica- End set from NCI
tion ($800$ images)
[19]
Yan et al. [12] 2-way End-to- ResNet18 Private data- $95.70$ $74.60$ $85.50$
(2021) classifica- End set from SRRS
tion Hospital
($1400$ images)
[12]
Yuan et al. [13] 2-way End-to- ResNet50 Private data $82.62$ $85.38$ $84.10$
(2021) classifica- End set ($22330$
tion images) [13]
Cho et al. [29] 5-way End-to- Inception- Private data- $51.70$
(2023) classifica- End ResNet-V2, set ($791$)
tion ResNet152 images [29]
Dash et al. [18] 3-way Hybrid Inception- SVM Public data- $90.62$ $81.24$ $81.24$
(2023) classifica- ResNet-V2 set from IARC
tion $292$ images
[18]
3 Methodology
In this section, we discuss two components of our proposed CAD system, which include the feature processing technique and the classifier. The complete setup of our proposed CAD system is shown in Figure 1, and different components are discussed below.
Figure 1: Flow diagram of proposed CAD system
3.1 Feature processing technique
This subsection has three parts. First, we extract features from three pre-trained deep learning descriptors of the ResNet family individually and concatenate the obtained features as well. Second, we extract features at the end of each convolution block of these ResNets and then again concatenate them. Third, we normalize the resulting feature set. The three components mentioned above are discussed below, which also give the motivation for making these choices.
3.1.1 Feature extraction using ResNets
ResNets have been recently shown to work well as pretrained deep learning descriptors in the field of medical imaging [11, 14]. These descriptors also solve the vanishing gradient problem common in deep neural networks. ResNet50 contains 50 layers in total, which includes 49 convolutional layers that are arranged in 16 residual blocks, as shown in Figure 2. It is a relatively lightweight model that makes it faster, and it captures simpler patterns. ResNet101 has 101 layers, which include 100 convolutional layers that are organized into 33 residual blocks, as depicted in Figure 3. It balances depth and complexity, which makes it suitable for tasks where more advanced patterns are to be identified. ResNet152 is the deepest of the three, with 152 layers, including 151 convolutional layers divided into 50 residual blocks, as illustrated in Figure 4. It is capable of capturing the most complex patterns.
Besides individually extracting features from these three ResNets, we concatenate their features as well to improve the prediction results. Next, we improve these descriptors, which is discussed next.
Figure 2: Architecture of ResNet50 descriptor
Figure 3: Architecture of ResNet101 descriptor
Figure 4: Architecture of ResNet152 descriptor
3.1.2 Feature extraction using CBF-ResNets
It has been shown in pretrained neural networks that detailed features like edges, color, and texture are captured by earlier convolution blocks while abstract features like shape and object, are captured by later convolution blocks [31, 17, 16]. Subsequently, these abstract features dominate the output of the network. Since detailed features of colposcopy cancer image classification are as important as abstract features, extraction of features at the end of each convolution block becomes necessary. Hence, for each ResNet, we extract all such features while cascading through the network. In general, ResNets have five convolution blocks. As discussed above, the first two blocks capture the detailed features, and the last three blocks capture the abstract features.
Directly combining features obtained at the end of each convolution block would dramatically increase the number of features, making the network both less efficient (due to the size of the feature set) and less effective (due to the contribution of the less relevant features). Hence, to address this, we need to perform feature selection.
Feature selection can be performed in many ways. In deep learning, one of the common approaches is pooling, which can be done using various methods such as Global Sum Pooling (GSP), Global Max Pooling (GMP), and Global Average Pooling (GAP) [20, 21]. As discussed earlier, GMP is effective in retaining the strongest activations by filtering out the noise. The initial two convolution blocks that provide detailed features also carry background noise. Hence, to eliminate this noise, we pass the output of the first two convolution blocks through GMP layers.
The GAP layer averages out the features over the whole image and does not filter out any background noise. The later three convolution blocks that provide abstract features do not have any noise from the background. Hence, we pass the output of the last three convolution blocks through GAP layers [32, 22]. This novel ensemble pooling strategy enhances both feature relevance and model efficiency. The final feature vector is formed by fusing the GMP and GAP feature vectors obtained from each convolution block.
We refer to our new descriptors as Cascaded Block Fused ResNets (CBF-ResNets). This includes CBF-ResNet50, CBF-ResNet101, and CBF-ResNet152, corresponding to ResNet50, ResNet101, and ResNet152, respectively. Eventually, features from the three new descriptors are concatenated. The complete step-by-step process, including feature extraction from each convolution block, feature selection using the GMP and GAP layers, fusion of features across all blocks, and final concatenation of features from multiple ResNet models, is outlined in Algorithm 1. Furthermore, we combine the extracted features from all the proposed CBF-ResNets (CBF-ResNet50, CBF-ResNet101, and CBF-ResNet152).
Algorithm 1 Cascaded Block Fused ResNet (CBF-ResNet) Feature Extraction and Concatenation
1:Input: Colposcopy cancer image dataset $\mathbf{D}$, containing $N$ images.
2:Output: Concatenated feature vector $\mathbf{F}_{\text{CBF}}$ from all three ResNet descriptors.
3:Definitions:
4:$\mathcal{I}\in\mathbf{D}$: A single colposcopy cancer image.
5:$r\in\{50,101,152\}$: ResNet model identifier (ResNet50, ResNet101, ResNet152).
6:$k\in\{1,2,3,4,5\}$: Index of convolutional blocks in the ResNet model.
7:$H_{k},W_{k},C_{k}$: Height, width, and number of channels of the feature maps from the $k^{th}$ convolution block, where $H_{k}$, $W_{k}$, and $C_{k}$ are positive integers.
8:$F_{k}\in\mathbb{R}^{H_{k}\times W_{k}\times C_{k}}$: Feature map extracted from the $k^{th}$ convolution block.
9:$\mathbf{G}_{k}^{M}\in\mathbb{R}^{C_{k}}$: GMP feature vector obtained from $F_{k}$.
10:$\mathbf{G}_{k}^{A}\in\mathbb{R}^{C_{k}}$: GAP feature vector obtained from $F_{k}$.
11:$\mathbf{F}_{\text{CBF}}^{(r)}\in\mathbb{R}^{d_{r}}$: Feature vector of ResNet-$r$ obtained by concatenating pooled features from all blocks, where $d_{r}$ is the dimensionality of the concatenated feature vector.
12:$\mathbf{F}\in\mathbb{R}^{d_{\text{CBF}}}$: Final concatenated feature vector from ResNet50, ResNet101, and ResNet152.
13:Feature Extraction:
14:for each image $\mathcal{I}\in\mathbf{D}$ do
15:    for each ResNet model $r\in\{50,101,152\}$ do
16:         for $k=1$ to $5$ (each convolution block) do
17:             Extract feature maps $F_{k}\in\mathbb{R}^{H_{k}\times W_{k}\times C_{k}}$ from the $k^{th}$ convolution block.
18:             if $k\in\{1,2\}$ then
19:                 Apply Global Maximum Pooling (GMP) to obtain $\mathbf{G}_{k}^{M}$:
20:
$\mathbf{G}_{k}^{M}=\left[\max_{(i,j)\in H_{k}\times W_{k}}F_{k}^{(c)}[i,j]\right]_{c=1}^{C_{k}}.$
21:             else
22:                 Apply Global Average Pooling (GAP) to obtain $\mathbf{G}_{k}^{A}$:
23:
$\mathbf{G}_{k}^{A}=\left[\frac{1}{H_{k}W_{k}}\sum_{i=1}^{H_{k}}\sum_{j=1}^{W_{k}}F_{k}^{(c)}[i,j]\right]_{c=1}^{C_{k}}.$
24:             end if
25:         end for
26:         Fuse pooled feature vectors for ResNet-$r$:
27:
$\mathbf{F}_{\text{CBF}}^{(r)}=\left[\mathbf{G}_{1}^{M}\mathbf{G}_{2}^{M}\mathbf{G}_{3}^{A}\mathbf{G}_{4}^{A}\mathbf{G}_{5}^{A}\right].$
28:    end for
29:    Concatenate feature vectors from all three ResNet models:
30:
$\mathbf{F}=\left[\mathbf{F}_{\text{CBF}}^{(50)}\mathbf{F}_{\text{CBF}}^{(101)}\mathbf{F}_{\text{CBF}}^{(152)}\right].$
31:end for
Next, the data is split into training and testing parts (coming from cross-validation, which is discussed in the Results section). Subsequently, features are normalized, as discussed below.
3.1.3 Feature normalization
We normalize the features to prevent the classification model from biased learning and ensure fair comparisons between different units or measurement scales. There are various ways available to normalize the features, i.e., z-scored normalization, variable stability scaling, min-max normalization, etc. [33]. For us, the min-max normalization technique, which scales the features in the range between 0 and 1, works the best. It is formulated as follows:
$s_{\text{scaled}}=\frac{s-s_{\text{min}}}{s_{\text{max}}-s_{\text{min}}},$ (1)
where $s_{\text{scaled}}$ is the scaled value of a feature, $s$ is the original value of the feature, $s_{\text{min}}$ is the minimum among all the feature values, and $s_{\text{max}}$ is the maximum value, again, among all the feature values. Next, the data is fed to the classifier, which is discussed below.
3.2 Classifier
SVM is a supervised learning method used to classify data. It works by finding a function to capture the relationships between different points, which is eventually used to separate them into classes [34]. SVM has emerged as a powerful tool for machine learning and pattern recognition tasks because of its ability to achieve high generalization performance, even with limited training data. As mentioned above, for the types and the CIN score classification, we have three classes. The standard SVM performs a two-class classification, which we discuss below. This can be easily generalized to a three-class classification by using a combination of classes.
We use linear SVM to classify the colposcopy images. This finds a linear hyperplane to separate the data points of the two classes and is formulated as follows [34]:
$\mathbf{f(x)}=\mathbf{w}^{T}\mathbf{x}+b,$ (2)
where $\mathbf{w}$ is the weight vector and $b$ is the bias. The two together are also known as the hyperplane parameters.
The optimal hyperplane is determined by maximizing the margin $\gamma=\frac{2}{\|\mathbf{w}\|}$ between the two classes. To achieve this, we minimize the value of $\|\mathbf{w}\|$, which leads to the following optimization problem:
$\min_{\mathbf{w}}\frac{1}{2}\|\mathbf{w}\|^{2},$ (3)
subject to the constraints
$y_{i}\left(\mathbf{w}^{T}\mathbf{x}_{i}+b\right)\geq 1,\quad\text{for}\ i=1,2,\dots,m,$ (4)
where $y_{i}$ is the label of $i^{th}$ training sample and $m$ are the number of samples.
4 Data preprocessing and experimental results
Here, we first describe the data preprocessing, and then give our experimentation results.
4.1 Data preprocessing
First, we discuss the data preprocessing for the International Agency for Research on Cancer (IARC) dataset, provided by the World Health Organization (WHO) [27]  [^2] This dataset is not transferrable and is meant for this research only. If anyone needs it further, then they may contact the main source.. As discussed in literature, this dataset is categorized based on the transformation zone (TZ). The TZ refers to the region around the cervical opening where the endocervix (inner part) and ectocervix (outer part) converge (where the squamous epithelium cells replace the columnar epithelium cells). The transformation zone is typically classified into three types, namely, Type1, Type2, and Type3. For this dataset, we follow the exact same preprocessing methodology outlined in [18].
The original colposcopy images are of $800\times 600$ pixels. We rescale all images to $224\times 224$ to optimize the computational resources. These images not only contain the cervical region (Region of Interest or ROI), but they also contain some artifacts that are not relevant. Therefore, we segment the images, which reduces the impact of additional artifacts beyond ROI. Some images are discarded during the segmentation process because the cervix region is not properly visible in the original images, and after segmentation, we are left with only irrelevant regions. Figure 5 shows the sample of original images and their segmented counterparts.
The total number of original images and the segmented images for the types classification are given in columns 2 and 3 of Table 2. As evident from this table, the data is imbalanced. To balance the types, we increase the number of images for the class with the fewest samples by a factor of 5 via rotation and flip. We match the images of the other two classes with the final number obtained above (again by rotating and flipping them). This results in 315 images for each class, as given in column 4 of Table 2. To further increase the data five times, we perform random variations of contrast, brightness, rotation, and translation. This results in 1575 images for each class for a total of 4725 images, as given in columns 5 and 6 of Table 2, respectively.
The colposcopy images given in the AnnoCerv dataset are categorized into three categories based on their scores, ranging from $0-10$. The images with scores $0-4$ fall into the first category (low-grade/CIN1), $5-6$ fall into the second category (high-grade/non-invasive cancer/CIN2+), and $7-10$ belong to the third category (high-grade/suspected invasive cancer/CIN2+) [28].
Next, we follow the same preprocessing steps as above for our second dataset. The results of the segmentation and the augmentation process are listed in Table 3.
Figure 5: Rows show the sample of original images and their segmented counterparts [27], Columns show; (a) Normal (b) Abnormal (c) Type 1 (d) Type 2 (e) Type 3 images
Table 2: Dataset balancing and augmentation for the IARC dataset
Category Original Segmented Rot. Rand. Var. No. of
Image Images & Flip (Cont., Bright., Images
Rot., Trans.)
Type1 318 226 315 1575 4725
Type2 106 63 315 1575
Type3 147 87 315 1575
  • Rand. Var. = Random Variation, Bright. = Brightness, Rot. = Rotation, Cont. = Contrast, and Trans. = Translation
Table 3: Dataset balancing and augmentation for the AnnoCerv dataset
Category Original Segmented Rot. Rand. Var. No. of
Score Image Images & Flip (Cont., Bright., Images
Rot., Trans.)
0-4 311 311 311 1555 4665
5-6 124 124 311 1555
7-10 96 96 311 1555
4.2 Experimental results
For training and testing, we use 5-fold and 10-fold cross-validation methods [35, 36]. In 5-fold, the data is split into roughly five equal parts, known as folds. Now, the model is trained and tested five times, with a different fold used as the test set in each round, while the other four folds are used for training. After all five rounds, the results are averaged to give a more reliable and unbiased measure of the model’s performance on the new data. Similarly, in 10-fold cross-validation, the process is the same, but the data is split into ten folds, and the model is trained and tested across ten iterations. To assess the performance of our system, we use standard metrics such as sensitivity, specificity, and accuracy.
Sensitivity measures the ability of the model to identify positive instances out of the total actual positives. It is calculated as follows:
$Sensitivity=\frac{TP}{TP+FN},$ (5)
where TP and FN mean True Positive and False Negative, respectively.
Specificity assesses the capability of the model to identify negative instances out of the total actual negatives. It is formulated as follows:
$Specificity=\frac{TN}{TN+FP},$ (6)
where TN and FP mean True Negative and False Positive, respectively.
Precision refers to the proportion of correctly predicted positive instances out of all instances that were predicted as positive. It is computed as follows:
$Precision=\frac{TP}{TP+FP},$ (7)
F1-score is the harmonic mean of sensitivity and precision, offering a balanced measure when both false positives and false negatives are important. It is calculated as follows:
$F1\text{-}score=\frac{2\times Precision\times Sensitivity}{Precision+Sensitivity}.$ (8)
Accuracy represents the overall correctness of the classification model and is calculated as the ratio of the correctly predicted instances to the total instances:
$Accuracy=\frac{TP+TN}{TP+FN+TN+FP}.$ (9)
The results for the types classification using 5-fold and 10-fold cross-validation on the IARC dataset are presented in Table 4 and Table 5, respectively. Here, the first column lists the feature extraction techniques, and the rest of the columns list the values of the performance metrics. The corresponding bar plots are given in Figure 6 and Figure 7, respectively. It is apparent from the tables and the figures that individual ResNet descriptors exhibit modest performance with an average of $90.89\%$), while a notable improvement is obtained by proposed CBF-ResNets descriptors with 93.97%.
Furthermore, when we combine and normalize the features of the original ResNet50, ResNet101, and ResNet152, the average performance improves to $96.06\%$. In contrast, the normalized combination of our CBF-ResNet50, CBF-ResNet101, and CBF-ResNet152 achieves a superior average performance of 97.92%. We compute this average performance by taking the mean of all five performance metrics obtained from both 5-fold and 10-fold cross-validation. We also compare our results with the 10-fold cross-validation results reported in [18]. Our method shows a substantial improvement over theirs, with gains of $10\%$ in precision, $16\%$ in F1-score, $8\%$ in specificity, $16\%$ in sensitivity, and $16\%$ in accuracy.
Table 4: Performance comparison of the proposed model using 5-fold cross-validation for the types classification on IARC dataset
Features Extraction Techniques Pre.(%) F1-score(%) Spec. (%) Sen. (%) Acc. (%)
ResNet50 89.48 89.45 93.43 88.97 91.93
CBF-ResNet50 92.31 92.25 94.26 92.12 93.54
ResNet101 89.42 89.41 93.99 88.70 92.23
CBF-ResNet101 93.87 93.83 96.40 93.20 95.34
ResNet152 89.52 89.49 93.53 89.29 92.10
CBF-ResNet152 93.39 93.38 95.68 92.25 94.53
ResNet50 + ResNet101 + 95.96 95.91 96.90 96.44 96.74
ResNet152 (Normalized)
CBF-ResNet50 + CBF-ResNet101 + 97.97 97.98 98.89 97.33 98.37
CBF-ResNet152 (Normalized)
Figure 6: Performance bar plot for the types classification with 5-fold cross-validation on IARC dataset
Table 5: Performance comparison of the proposed model using 10-fold cross-validation for the types classification on IARC dataset
Features Extraction Techniques Pre.(%) F1-score(%) Spec. (%) Sen. (%) Acc. (%)
ResNet50 89.74 89.65 93.06 89.27 91.80
CBF-ResNet50 93.34 93.31 95.70 92.38 94.60
ResNet101 90.07 90.02 93.76 89.22 92.25
CBF-ResNet101 94.22 94.19 96.35 93.56 95.40
ResNet152 90.43 90.41 94.43 89.38 92.76
CBF-ResNet152 93.68 93.70 95.78 92.15 94.56
ResNet50 + ResNet101 + 95.31 95.27 96.04 96.04 96.02
ResNet152 (Normalized)
CBF-ResNet50 + CBF-ResNet101 + 97.63 97.60 98.11 97.48 97.88
CBF-ResNet152 (Normalized)
Figure 7: Performance bar plot for the types classification with 10-fold cross-validation on IARC dataset
Next, the results for the CIN score classification on the AnnoCerv dataset using 5-fold and 10-fold cross-validation are given in Table 6 and Table 7, respectively. The respective bar plots are given in Figure 8 and Figure 9. This follows the same pattern as the types classification. Here, the individual ResNets descriptors achieve an average performance of $87.97\%$, while our individual CBF-ResNet descriptors improve the performance, reaching 89.41%.
The normalized combination of ResNet50, ResNet101, and ResNet152 further enhances the performance, achieving an average of $91.23\%$. A more significant gain is achieved by combining and normalizing the features extracted from our proposed CBF-ResNet50, CBF-ResNet101, and CBF-ResNet152 descriptors, resulting in an average performance of 92.97%. The average performance is computed in the same manner as described for the IARC dataset. Currently, no competitive work is available on this dataset for comparison.
Table 6: Performance comparison of the proposed model for the CIN score classification with 5-fold cross-validation on AnnoCerv dataset
Features Extraction Techniques Pre.(%) F1-score(%) Spec. (%) Sen. (%) Acc. (%)
ResNet50 87.95 87.86 91.49 86.56 89.83
CBF-ResNet50 89.29 89.20 92.25 87.34 90.61
ResNet101 86.49 86.37 90.20 85.87 88.76
CBF-ResNet101 89.48 89.40 92.28 87.71 90.73
ResNet152 87.27 87.21 91.57 85.65 89.60
CBF-ResNet152 89.12 89.02 91.31 88.30 90.31
ResNet50 + ResNet101 + 91.22 90.92 92.17 92.24 92.17
ResNet152 (Normalized)
CBF-ResNet50 + CBF-ResNet101 + 93.28 93.18 92.39 94.14 93.54
CBF-ResNet152 (Normalized)
Figure 8: Performance bar plot for the CIN score classification with 5-fold cross-validation on AnnoCerv dataset
Table 7: Performance comparison of the proposed model for the CIN score classification with 10-fold cross-validation on AnnoCerv dataset
Features Extraction Techniques Pre.(%) F1-score(%) Spec. (%) Sen. (%) Acc. (%)
ResNet50 87.40 87.28 90.63 86.35 89.17
CBF-ResNet50 88.97 88.91 91.68 87.19 90.18
ResNet101 86.17 86.10 90.55 84.24 88.44
CBF-ResNet101 88.58 88.48 91.28 86.31 89.64
ResNet152 87.00 86.94 91.21 85.62 89.34
CBF-ResNet152 88.73 88.62 91.40 86.41 89.71
ResNet50 + ResNet101 + 90.23 89.49 90.27 92.63 91.03
ResNet152 (Normalized)
CBF-ResNet50 + CBF-ResNet101 + 92.72 92.45 92.41 93.08 92.66
CBF-ResNet152 (Normalized)
Figure 9: Performance bar plot for the CIN score classification with 10-fold cross-validation on AnnoCerv dataset
5 Explainable AI (XAI)
XAI refers to the process of understanding and explaining the decisions of machine learning or deep learning models. This is especially important in the deep learning context due to the inherent complexity of these models and the challenges involved in interpreting their decision-making processes. There are several XAI techniques available to explain the decision of these models, i.e., CAM (Class Activation Mapping), Grad-CAM (Gradient-weighted Class Activation Mapping), SHAP (Shapley Additive Explanations), and LIME (Local Interpretable Model-agnostic Explanations), etc. [37, 38].
Among the previously mentioned XAI techniques, we adopt SHAP and LIME due to their model-agnostic nature. However, applying these methods individually to our model does not provide the level of explanation we require. Hence, we propose a novel ensemble approach that combines the strengths of both SHAP and LIME.
Among the two available datasets, the IARC dataset is used for the types classification, while the AnnoCerv dataset is used for the CIN score classification. Performing explainability on the types and the CIN score is difficult, as we cannot verify the type and the CIN score with naked eyes. Hence, the easiest way to verify the explainability results is to apply explainability to the abnormal images, which are also available in the IARC dataset. Due to computational constraints, we work with about one-third of such abnormal images [39].
It is a known fact that cervical cancer typically occurs in the region around the cervix [40]. Hence, for the original abnormal images being studied, we identify the abnormality and mark this region with a digital pen. For the SHAP model, the output is in the form of two sets of images, the SHAP normal image and the SHAP abnormal image, with the region contributing to the decision-making of normal-abnormal highlighted in red color. If the marked cancerous region in the original abnormal image matches with the red region in the SHAP abnormal image, then we say it is a correct explanation. In all the other cases, we say the explanation is incorrect.
For the LIME technique, the output is only in the form of one image where the region that has contributed to the decision-making of normal-abnormal is present in the image, and the region that has not contributed is blackened. If the marked cancerous region in the original abnormal image matches with the un-blackened region in the LIME image, then we say that is a correct explanation. In all the other cases, we say the explanation is incorrect.
The results of this approach for about one-third of available abnormal images ($122$) are presented in Table 8. As shown in the table, the first row highlights the performance of the SHAP technique, which independently provides correct explanations for $96$ images and incorrect explanations for $26$ images, achieving a performance of $78.68\%$. The second row presents the performance of LIME, which independently generates the correct explanations for $92$ images and incorrect explanations for $30$ images, achieving a performance of $75.40\%$.
Table 8: Comparison of the XAI techniques for the abnormal images
XAI No. of Correctly No. of Incorrectly Performance
Techniques Explained Images Explained Images (%)
Independent SHAP 96 26 78.68
Independent LIME 92 30 75.40
Collective Performance $\mathrm{I}$ 117 5 95.90
Collective Performance $\mathrm{II}$ 119 3 97.54
Figure 10: Correctly explained images by ensemble of SHAP and LIME techniques on selection images from [27].
The third row of the table lists the data for the collective performance $\mathrm{I}$. Here, initially, LIME is applied, and then for those images where LIME fails, SHAP is applied. Thus, we achieve correct explanation for $117$ images ($92$ images by LIME and $25$ images by SHAP) and incorrect for $5$ images, achieving a performance of $95.90\%$. Similarly, the fourth row of the table lists the data for the collective performance $\mathrm{II}$. Here, initially, SHAP is applied, and then, for those images where SHAP fails, LIME is applied. Again, we achieve correct explanation for $119$ images ($96$ images by SHAP and $23$ images by LIME) and incorrect for $3$ images, achieving a performance of $97.54\%$.
A few abnormal images that are correctly explained by both SHAP and LIME are given in Figure 10. Since the region contributing to the decision in our classification model is correctly identified by our two ensemble XAI techniques in most of the cases, we can confidently say that our classification model performs correct classification.
6 Conclusions and future work
We propose a CAD system for better cervical cancer image classification. Due to their recent success in medical imaging and their ability to avoid the vanishing gradient problem, ResNets are preferred here as feature extraction descriptors. Since lower and higher numbered ResNets capture simple and complex patterns, respectively, we use three such descriptors, i.e., ResNet50, ResNet101, ResNet152, which is our first model.
Different convolution blocks of pretrained neural networks give different features. Detailed features are available in the earlier blocks, and abstract features are available in the later blocks. These abstract features dominate the final output of the network, while detailed features are equally important for cervical cancer image classification. Hence, we improve the above-listed ResNets by extracting features at the end of each convolution block in a cascading manner. None of the earlier studies in cervical cancer classification have done this. Hence, this is innovative here. After extracting features from each convolution block, we apply a novel feature selection technique, where Global Max Pooling is used for detailed features, and Global Average Pooling is used for abstract features. This process results in three new descriptors, termed as CBF-ResNet50, CBF-ResNet101, and CBF-ResNet152. These constitute our second set of models.
It has been shown that combining features from different pretrained CNNs can improve performance. Therefore, we combine and normalize the features extracted from our first set of models (i.e., the three standard ResNets) and our proposed Cascaded Block Fused ResNets, which constitute our third and fourth sets of models, respectively. A linear SVM is then used for classification.
We perform our experiments on two datasets, namely, the IARC dataset provided by WHO and the AnnoCerv dataset. On the IARC dataset, the standard ResNets achieves an average performance of $90.89\%$, while our best model achieves a significantly higher average performance of 97.92%. Compared to the best available competitive approach on this dataset, our method shows an average improvement of 13.20%. On the AnnoCerv dataset, the standard ResNets attain an average performance of $87.97\%$, whereas our best model achieves an improved average performance of 92.97%. To the best of our knowledge, there is currently no competitive approach available for comparison on this dataset.
We also study the explainability of our classification models where we propose a new ensemble of SHAP and LIME XAI techniques. It is a well-known fact that cervical cancer originates around the cervix region. The region that contributes to the decision-making in our classification models, as identified by our ensemble of XAI techniques, turns out to be around the cervix in 97% of the cases. Hence, this validates the correctness of our classification models.
One of the future work directions is to test our CAD system on a much larger dataset, as WHO has mentioned that they would provide such a dataset soon. A second direction for future work is to study the implicit relation between different components [41]. Another future work could be to formulate the combination of different deep neural networks as an optimization problem as done in different domains [42, 43]. Finally, it would be good to explore the possibility of approximate computing [44, 45] in this context.
[…]
*Acknowledgements We would like to thank Mr. Siddartha Chennareddy and Mr. Karthik Boddupalli for their valuable discussions on various aspects of this research.
[…]
*Authors Contributions Saurabh Saini performed investigation, data curation, and writing - original draft. Kapil Ahuja conceptualised, performed project administration, and writing - review & editing. Akshat S. Chauhan performed formal analysis and software preparation.
[…]
*Funding This study was not funded. The authors have no relevant financial or non-financial interests to disclose.
[…]
*Data availability The IARC colposcopy dataset can be obtained upon request [27], while the AnnoCerv dataset is available in [28].
Declarations
[…]
*Competing Interest The authors state they have no financial or personal conflicts related to this research, authorship, or publication.
[…]
*Ethical Approval No experiments involving human participants or animals were conducted by the authors in this study.